OpenStudy (anonymous):

The rate at which a plant grows is constant. At the 3rd week it is 12 inches tall and at the 5th week it is 20 inches tall. If this information is plotted on a coordinate plane, what is the y-coordinate of the the point (1, y)? A) 1 B) 2 C) 4 D) 9

OpenStudy (anonymous):

@sleepyjess can u help me please?

OpenStudy (anonymous):

yea looking hold on

OpenStudy (anonymous):

ok I think is d but not sure

OpenStudy (danjs):

Y = k* x Use one of the points to find the constant k

OpenStudy (anonymous):

its c

OpenStudy (danjs):

3rd week 12 inches (,2,12) 12 = k *3 k=4

OpenStudy (danjs):

y = 4x

OpenStudy (anonymous):

12+4+4=20

OpenStudy (danjs):

x = 1, y = 4(1) = 4

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

its C

OpenStudy (anonymous):

thxs:)

OpenStudy (anonymous):

np anytime

OpenStudy (anonymous):

need anymore